OpenStudy (anonymous):

evaluate the following: integral sign 3/ x(x+3) (dx)

OpenStudy (blockcolder):

Use the technique of partial fractions here.

OpenStudy (anonymous):

somebody help me with this integration problems: \[\int\limits_{0}^{pi/2}\sin^4 x dx\] \[\int\limits_{1}^{e}\ln x dx\] \[\int\limits_{1}^{3}x^3 dx\] \[\int\limits_{}dy/(4y^2-24y+27)^3/2\]

OpenStudy (blockcolder):

1. Use the identity sin^2(x)=(1-cos(2x))/2 twice. 2. Use integration by parts. 3. Use the formula \[\int x^n dx=\frac{x^{n+1}}{n+1}\] 4. I can't understand the question.

OpenStudy (anonymous):

which part? the 3/2 is the power can u please show me the complete solution of them all?

OpenStudy (anonymous):

@blockcolder please do help me ....

OpenStudy (blockcolder):

\[\begin{align} \int \frac{3}{x(x+3)} dx&=\int \frac{1}{x}-\frac{1}{x+3}\ dx\\ &=\ln{x}+\ln(x+3)+C\\ &=\ln{\left (\frac{x}{x+3} \right )}+C \end{align}\]

OpenStudy (anonymous):

thank you so much your a genius

OpenStudy (blockcolder):

\[\begin{align} \int_{0}^{\pi/2} \sin^4x dx&=\int_{0}^{\pi/2} \left [ \frac{1+\cos(2x)}{2} \right ]^2 dx\\ &=\frac{1}{4} \int_{0}^{\pi/2} 1+2\cos(2x)+\cos^2(2x) dx\\ &=\frac{1}{4} \int_{0}^{\pi/2} 1+2\cos(2x)+\frac{1}{2}(1+\cos(2x))\ dx\\ &=\frac{1}{4} \left [ x+\sin(2x)+ \frac{x}{2}+\frac{\sin(2x)}{4} \right ]|_{0}^{\pi/4}\\ &= \frac{3\pi}{16} \end{align}\]

OpenStudy (anonymous):

thank you so much @blockcolder

OpenStudy (anonymous):

@blockcolder how about the others?

OpenStudy (blockcolder):

\[\int_{1}^{e} \ln{x}\ dx=(x \ln{x} -x)|_{1}^{e}=1\]

OpenStudy (blockcolder):

\[\Large \int_{1}^{3}x^3\ dx=\frac{1}{4}x^4|_1^3=20\]

OpenStudy (anonymous):

this is right wow ...

OpenStudy (anonymous):

@blockcolder how about the last one.

OpenStudy (blockcolder):

Click "Show steps" next to the answer. It's too long to be typed here. http://www.wolframalpha.com/input/?i=integral+of+1%2F%284x^2-24x%2B27%29^%283%2F2%29